Broward Manufacturing recently reported the following information:

Net income $777,000
ROA 8%
Interest expense $264,180
Accounts payable and accruals $950,000

Broward's tax rate is 25%. Broward finances with only debt and common equity, so it has no preferred stock. 40% of its total invested capital is debt, and 60% of its total invested capital is common equity. Calculate its basic earning power (BEP), its return on equity (ROE), and its return on invested capital (ROIC). Do not round intermediate calculations. Round your answers to two decimal places.

BEP:   %

ROE:   %

ROIC:   %
Answer #1

Net Income = 777000

ROA = Net Income/Total Assets

Hence, total assets = 777000/8%

= $9,712,500

EBIT = 777000/(1-25%)+ 264180

= $1,300,180

Basic Earning power = EBIT/Total assets

= 1300180/9712500

= 13.38666%

i.e. 13.39%

Equity = 9712500*60% = $5827500

ROE = Net Income/Equity

= 777000/5827500

= 13.3333%

ROIC = EBIT(1-Tax)/Total Invested capital

= 1300180(1-25%)/9712500

= 10.04%